what is the standard form equation of the conic given ? 4x^2-5y^2-16x-30y-9= 0
Answer by stanbon(75887)

what is the standard form equation of the conic given ? 4x^2-5y^2-16x-30y-9= 0
--------------
Complete the square for the x-terms and y-terms respectively:
4(x^2-4x+4) -5(y^2+6y+9)=9+16-45
4(x-2)^2 - 5(y+3)^2 = -20
-(x-2)^2/5 + (y+3)^2/4 = 1
========
Cheers,
Stan H.